OpenStudy (anonymous):

Which of the following equations describe the line shown below? All that apply. CORDS (-6,2) and (-1,-4) A. y + 4 = -(x + 1) B. y - 2 = -(x + 6) C. y - 2 = -(x + 6) D. y + 4 = -(x + 1) E. y - 4 = -(x - 1) F. y - 6 = -(x + 2)

OpenStudy (paxpolaris):

none of them

OpenStudy (anonymous):

there are alteast two of them that are correct.

OpenStudy (paxpolaris):

either you're writing something wrong or it's none of them..

OpenStudy (anonymous):

i am literally copy pasting.

OpenStudy (paxpolaris):

it should be: \[a.\ \ \Large y+4=-\frac65(x+1)\] and it should be: \[c.\ \ \Large y-2=-\frac65(x+6)\] and

OpenStudy (anonymous):

actually i just saw what is wrong

OpenStudy (anonymous):

whenever i copy paste it it's not doing the fraction!

OpenStudy (anonymous):

derp

OpenStudy (anonymous):

your answers were correct.

OpenStudy (paxpolaris):

if you get the slope 'm' right: you can write the equation as: \[\huge y-k=m \left( x-h \right)\] for any point (h,k) on the line